# GOT4Rec: Graph of Thoughts for Sequential Recommendation

### v1

@@ Abstract
With the advancement of large language models (LLMs), researchers have explored various methods to optimally leverage their comprehension and generation capabilities in sequential recommendation scenarios. However, several challenges persist in this endeavor. Firstly, most existing approaches rely on the input-output prompting paradigm, which can result in irrelevant or inaccurate responses. Secondly, while there have been attempts to enhance LLMs using prompting strategies such as chain-of-thought (CoT), these efforts have not fully harnessed the reasoning abilities of LLMs or effectively captured the multifaceted information contained within user sequences. To address these limitations, we propose GOT4Rec, a sequential recommendation method that utilizes the graph of thoughts (GoT) prompting strategy. Specifically, we identify and utilize three key types of information within user history sequences: short-term interests, long-term interests and collaborative information from other users. Our approach enables LLMs to independently reason and generate recommendations based on these distinct types of information, subsequently aggregating the results within the GoT framework to derive the final recommended items. This method allows LLMs, with enhanced reasoning capabilities, to more effectively consider the diverse information within user sequences, resulting in more accurate recommendations and more comprehensive explanations. Extensive experiments on real-world datasets demonstrate the effectiveness of GOT4Rec, indicating that it outperforms existing state-of-the-art baselines. Our code is available at https://anonymous.4open.science/r/GOT4Rec-ED99.
@@ 1 Introduction
Sequential recommendation has long been a significant research field, with numerous methods proposed to explore chronological dependencies within user sequences (Kang and McAuley 2018; Zhou et al. 2020). Despite considerable advancements, they remain constrained by the limited knowledge available from datasets. To overcome this, it is crucial to integrate real-world knowledge into sequential recommendation models, enabling them to more effectively comprehend and reason about preference patterns within user behavior sequences (Hou et al. 2022; Li et al. 2023a).
@@
Recently, large language models (LLMs) have garnered significant attention due to their impressive natural language comprehension and generation abilities. Consequently, numerous LLM-based sequential recommendation works (Xi et al. 2023; Sanner et al. 2023; Hou et al. 2024b) have emerged to exploit these capabilities, aiming to capture user interests from interaction sequences and leverage LLMs’ extensive real-world knowledge for recommendations. However, many of these approaches rely solely on the input-output (IO) prompting paradigm, underutilizing LLMs’ reasoning abilities. This limitation arises from a disconnect between the real-world knowledge embedded in LLMs and the specific needs of recommender systems, often resulting in task-irrelevant or inaccurate outcomes. To address this, several studies have incorporated advanced prompting strategies to enhance LLMs’ performance in sequential recommendation. Among these is SLIM (Wang et al. 2024), a knowledge distillation module which employs chain-of-thought (CoT) (Wei et al. 2022) prompting to enable step-by-step reasoning in sequential recommendation, transferring knowledge from a teacher model to a student model. However, it treats the user sequence as a monolithic entity and only taps into the elementary reasoning abilities of LLMs, insufficient for adequately reasoning through the diverse preference information (e.g., long-term and short-term interests, spatio-temporal factors) within user sequences.
@@
Overall, there are two major challenges that limit the effectiveness of LLMs in the sequential recommendation scenario. The first challenge is the difficulty in explicitly capturing various user preference information by merely prompting the behavior sequence. Traditional neural sequential recommenders have been actively extracting and fusing different types of information as features to predict the next item in the sequence. In contrast, current LLM-based sequential recommenders often perform minimal processing of the sequence itself, simply placing the entire sequence directly into the prompt. This approach can easily mislead the model, leading to inaccurate recommendations. For instance, in Figure 1, the items predicted by SLIM consist solely of snack bars, while the ground truth is a fruit nut mix product. The second challenge arises from the complexity introduced by incorporating multiple types of information, which transforms sequential recommendation into a complex reasoning task involving multiple sub-problems, necessitating LLMs with enhanced reasoning capabilities. Existing researches have demonstrated that simple reasoning approaches, such as IO and CoT, are insufficient for such tasks (Besta et al. 2024; Yang et al. 2024). In contrast, the graph of thoughts (GoT) method offers a more effective solution by decomposing the problem into more manageable components. Unlike CoT, which employs a single chain of thought, GoT models the reasoning process through networked reasoning. This approach allows GoT to break down complex reasoning tasks into smaller, more tractable sub-tasks, solve them individually, and then integrate the results to form a comprehensive solution. In Figure 1, we illustrate how GoT is employed to reason about three categories of products (i.e., “Probiotic Snack Bars”, “Healthy Snack Mixes” and “Low-Sugar, Low-Carb Cookies”) that the user might be interested in, enabling the LLM to independently recommend items within these categories and combine them in the end. In contrast, SLIM’s reasoning remains focused on snack bars.
@@
Figure 1: Comparison of the LLM output and predictions for the next item generated by SLIM (Wang et al. 2024) and our proposed GOT4Rec.
@@
To address the challenges mentioned above, this paper proposes GOT4Rec, a novel method that introduces the graph of thoughts (GoT) framework into the field of sequential recommendation. GOT4Rec optimally harnesses the reasoning capabilities of LLMs while effectively integrating multiple sources of information from user sequences. Specifically, we employ the GoT prompting strategy to extract three critical sources of information from user sequences: short-term interests, long-term interests, and collaborative interests from other users with similar preferences. Unlike previous methods that treat the sequence as a whole,our approach considers multiple aspects of information and better utilizes the reasoning abilities of LLMs, leading to more accurate and interpretable recommendations. Our key contributions can be summarized as follows:
@@
  • To the best of our knowledge, we are the first to apply the graph of thoughts framework within the field of sequential recommendation.
@@
  • We introduce GOT4Rec, a method that optimally leverages the reasoning capabilities of large language models (LLMs) to comprehensively capture and utilize the rich information embedded within user sequences.
@@
  • Extensive experiments conducted on three datasets demonstrate that our method outperforms traditional neural sequential models and other prompting strategies.
@@ 2 Related Work
2.1 Traditional Neural Sequential Recommenders
@@
Traditional neural sequential recommendation systems aim to capture sequential dependencies in user behavior sequences to model dynamic user preferences, as seen in methods like GRU4Rec (Hidasi et al. 2016). As deep learning develops, techniques like self-attention and graph neural networks have become foundational in sequential recommenders (Kang and McAuley 2018; Zhou et al. 2020; Wu et al. 2019; Zhang et al. 2022, 2023). These methods however, predominantly rely on sequence modeling capabilities and often inadequately incorporate textual information. Recently, research on transferable item representations (Hou et al. 2022; Li et al. 2023a) has gained attention. Despite their promise, these approaches remain constrained by limited datasets and fail to fully leverage real-world knowledge.
2.2 LLM-Based Sequential Recommenders
With the advent of LLMs, numerous research endeavors have explored leveraging their advanced language comprehension and generation abilities. LLMs can be integrated either as feature enhancers (Xi et al. 2023; Lin et al. 2024; Liu et al. 2024) or rankers/scorers (Dai et al. 2023; Harte et al. 2023; Li et al. 2023b; Sanner et al. 2023; Hou et al. 2024b). For example, ReLLa (Lin et al. 2024) employs semantic user behavior retrieval to improve data quality and introduces retrieval-enhanced instruction tuning to enhance few-shot recommendation. LLMRank (Hou et al. 2024b) formalizes the recommendation problem as a conditional ranking task and utilizes LLMs as zero-short rankers. Although these approaches are promising, existing research has not yet fully capitalized on the reasoning capabilities of LLMs.
2.3 LLM Prompting Strategies
@@
Numerous prompting approaches have been proposed to exploit the reasoning capabilities of LLMs. Chain-of-thought (CoT) (Wei et al. 2022) introduces intermediate reasoning steps to enhance LLM performance, while Chain of Thought with Self-Consistency (CoT-SC) (Wang et al. 2023) refines this by generating multiple CoTs and selecting the best output. Tree of thoughts (ToT) (Yao et al. 2024) and graph of thoughts (GoT) (Besta et al. 2024) further extend these methods by modeling reasoning as a tree or graph, respectively, to better generate and aggregate different thoughts. In the recommendation domain, SLIM (Wang et al. 2024) utilizes a CoT-based knowledge distillation module to transfer the step-by-step reasoning capabilities from a teacher model to a student model. However, its reliance on basic CoT prompting limits its ability to fully capture the abundant information within user sequences.
@@
Figure 2: An example of graph decomposition in the proposed GOT4Rec method. The user interaction sequence is divided to facilitate different types of reasoning: the last few items are used for short-term preference reasoning, while the entire sequence informs long-term preference reasoning. Additionally, sequences from similar users are retrieved for collaborative reasoning. In the thought graph, these distinct thoughts are generated and subsequently aggregated to capture and integrate the various aspects of information within the user interaction sequence.
@@ 3.1 Overview of Thought Graph
When interacting with LLMs, we input messages (prompts) and LLMs respond with generated outputs (thoughts). Building on the GoT framework (Besta et al. 2024), we model our GOT4Rec method as a tuple $(G,T)$, where $G$ represents the reasoning process and $T$ denotes the thought transformations. Specifically, we model the reasoning process as a directed graph $G=(V,E)$, where $V$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. Each vertex represents a thought, encapsulating a solution to the current recommendation step along with other relevant global information. A directed edge $(t_{1},t_{2})$ signifies a dependency between thoughts, indicating that thought $t_{2}$ is generated by LLMs based on $t_{1}$.
In our method, we employ the generation transformation $T_{G}$ and aggregation transformation $T_{A}$ from the GoT framework. The generation transformation generates one or more new thoughts based on an existing single thought $v$. In this operation, new vertices and edges are generated: $V_{+}=\{v_{1+},...,v_{k+}\}$ and $E_{+}=\{(v,v_{1+}),...,(v,v_{k+})\}$, where $v_{1+},...,v_{k+}$ are the new thoughts generated by $v$. Analogous reasoning steps such as CoT can be incorporated into this process. The aggregation transformation, on the other hand, combines multiple thoughts into a new, consolidated thought, reinforcing their strengths while mitigating their weaknesses. In this operation, a new vertex $v_{+}$ is created: $V_{+}=\{v_{+}\}$ and $E_{+}=\{(v_{1},v_{+}),...,(v_{k},v_{+})\}$, where $v_{1},...,v_{k}$ are the aggregated thoughts.
@@
3.2 Short-Term Reasoning Process
Numerous studies have emphasized the importance of understanding both a user’s dynamic short-term and stable long-term preferences in recommendation systems (Yu et al. 2019; Zheng et al. 2022). In our GOT4Rec method, we enable LLMs to extract these two types of preferences separately and then synergize the most relevant aspects of each.
To reason about short-term preferences, we select the last few interactions from the user sequence $S_{u}$ as $S_{short}$. We then apply and enhance the zero-shot CoT prompting strategy from (Wang et al. 2024) within the generation transformation $T_{G}$ to effectively capture the user’s short-term preferences and identify three categories that the user is likely to favor. As illustrated in Figure 3 as summarizing prompt, the generation transformation $T_{G}$ involves two key steps:
  • Step 1. Summarize user’s short-term preferences based on the given $S_{short}$.
@@ 3.2 Short-Term Reasoning Process
For each identified category, we prompt the LLMs to generate $N$ items the user is most likely interested in, repeating this process three times. The prompt template is provided in Figure 3 as recommendation prompt. After generating the item sets, we utilize the aggregation transformation $T_{A}$, where the three item sets undergo a voting process by the LLMs. This results in a new set of $N$ items that are most likely to interest the user within each category. After aggergation, we have three final item sets, which are subjected to a final round of voting by the LLMs to determine the top-$N$ items $I_{short}$ that best represent the user’s short-term preferences. The voting prompt is also provided in Figure 3.
3.3 Long-Term Reasoning Process
To capture long-term preferences, we extend the input to include the entire user interaction sequence. Similar to the short-term reasoning process, we employ the zero-short CoT prompting strategy within the generation transformation $T_{G}$, as shown in Figure 3. This allows LLMs to effectively differentiate and identify short-term and long-term preferences. After generating the relevant items, the final set of top-$N$ items $I_{long}$, representing the user’s long-term preferences, is determined through the aggregation transformation $T_{A}$.
@@ 3.4 Collaborative Reasoning Process
Collaborative filtering is widely used in various recommendation scenarios (Koren, Rendle, and Bell 2021), modeling users’ preference based on their past interactions. In our method, we leverage the all-mpnet-base-v2 (Reimers and Gurevych 2020), a sentence-transformers model that maps sentences to a vector space, to generate an embedding vector for the current user’s interaction sequence, allowing us to retrieve a set of sequences $S_{co}$ from other users most similar to the current user’s sequence. Details of this retrieval process are provided in the following section.
Once we have obtained the similar user sequences $S_{co}$, we employ the zero-short CoT prompting strategy to generate three sets of top-$N$ items that the current user may be interested in. The two-step generation prompts, illustrated in Figure 3 as collaboration prompt, are as follows:
  • Step 1. Summarize the shared preferences between the current user and other users based on the given $S_{co}$.
  • Step 2. Based on the summarized preferences from Step 1, recommend $N$ items selected from $S_{co}$ that the current user is likely to prefer.
The three sets of items generated through this process are then aggregated and selected by the LLMs using the aggregation transformation $T_{A}$. This results in the final set of top-$N$ items $I_{co}$ that best reflect the collaborative preferences of users with similar interests to the current user.
At this point, we have obtained three distinct sets of items: $I_{short}$, which reflects the user’s short-term preferences; $I_{long}$, which captures the user’s long-term preferences; and $I_{co}$, which represents the collaborative preferences derived from other users. Finally, using the aggregation transformation $T_{A}$, we enable LLMs to synthesize these multiple sources of information to generate the final set of top-$N$ recommended items $I_{fin}$. The prompt template for this final aggregation is the same as the voting prompt.

### v2

@@ Abstract
With their vast open-world knowledge and reasoning abilities, large language models (LLMs) have become a promising tool for sequential recommendation. Researchers have explored various methods to harness these capabilities, but most existing approaches rely on simple input-output prompting, failing to effectively bridge the gap between LLMs’ general knowledge and the specific needs of recommendation tasks. While reasoning strategies like chain-of-thought (CoT) have been introduced to enhance performance, they often produce inaccurate recommendations due to underutilized user preference information and insufficient reasoning depth. To address these challenges, we propose GOT4Rec, a novel sequential recommendation method leveraging the graph of thoughts (GoT) reasoning strategy. Our method focuses on three key types of information in user histories: short-term interests, long-term interests and collaborative information from other users. It enables LLMs to reason independently and generate recommendations, subsequently aggregating results to derive final items. This method allows LLMs, with enhanced reasoning capabilities, to better utilize the user sequence information, producing more accurate recommendations and comprehensive explanations. Extensive experiments on real-world datasets demonstrate the effectiveness of GOT4Rec, outperforming existing state-of-the-art baselines with an average improvement of 37.11%. Our code is available at https://anonymous.4open.science/r/GOT4Rec.
@@ 1 Introduction
Sequential recommendation has long been a significant research field, with numerous methods proposed to explore chronological dependencies within user sequences Kang and McAuley (2018); Zhou et al. (2020). Despite considerable advancements, they remain constrained by the limited knowledge available from datasets. To overcome this, it is crucial to integrate real-world knowledge into sequential recommendation models, enabling them to more effectively comprehend and reason about preference patterns within user behavior sequences Hou et al. (2022); Li et al. (2023a).
@@
Large language models (LLMs) have recently attracted significant attention for their expansive open-world knowledge and advanced reasoning abilities. Consequently, numerous LLM-based sequential recommendation methods Xi et al. (2024); Sanner et al. (2023); Hou et al. (2024b) have emerged, aiming to capture user interests from interaction sequences while leveraging LLMs’ comprehensive real-world knowledge for recommendations. Many of these approaches rely solely on the input-output prompting paradigm, which underutilizes the reasoning potential of LLMs. As a result, it often produces task-irrelevant or inaccurate outcomes. To address this problem, several studies have explored advanced reasoning strategies to boost LLMs’ performance in sequential recommendation. One example is SLIM Wang et al. (2024), a knowledge distillation module which employs chain-of-thought (CoT) Wei et al. (2022) prompting to enable step-by-step reasoning in sequential recommendation, transferring knowledge from a teacher model to a student model. However, it only utilizes the basic reasoning abilities of LLMs. This is insufficient for effectively processing the rich preference information embedded in user sequences, such as long-term and short-term interests or spatio-temporal patterns. Consequently, it often produces inaccurate recommendations. As illustrated in Figure 1, SLIM’s CoT-based reasoning focuses narrowly on predicting snack bars, while the true preference is a fruit nut mix, highlighting its limitations in handling diverse preference structures.
@@
Therefore, the challenge lies in integrating diverse user preferences into a cohesive reasoning framework, turning sequential recommendation into a multi-faceted task requiring advanced reasoning capabilities. Studies have shown that simple approaches like CoT are inadequate for tasks that demand decomposition into sub-tasks Besta et al. (2024); Yang et al. (2024). In contrast, the graph of thoughts (GoT) method offers a more effective solution by decomposing the problem into more manageable components. Unlike CoT, which employs a single chain of thought, GoT models the reasoning process through networked reasoning. This approach allows GoT to break down the complex user preference reasoning tasks into smaller, more tractable sub-tasks, solve them individually, and then integrate the results to form a comprehensive solution. As illustrated in Figure 1, GoT reasons over multiple product categories that the user might be interested in: Probiotic Snack Bars, Healthy Snack Mixes and Low-Sugar, Low-Carb Cookies, enabling the LLM to independently generate recommendations within each category and combine them to correctly predict the ground truth item: fruit nut mix. While SLIM’s recommendation remains focused on snack bars, showcasing its narrower approach.
@@
Figure 1: Comparison of the LLM output and predictions for the next item generated by SLIM and our proposed GOT4Rec.
@@
To address the challenges mentioned above, this paper proposes GOT4Rec, a novel method that introduces the graph of thoughts (GoT) framework into the field of sequential recommendation. GOT4Rec optimally harnesses the reasoning capabilities of LLMs while effectively integrating multiple sources of information from user sequences. Specifically, we employ the GoT reasoning strategy to extract three critical sources of information from user sequences: short-term interests, long-term interests, and collaborative interests from other users with similar preferences. Unlike previous methods that treat the sequence as a whole,our approach considers multiple aspects of information and better utilizes the reasoning abilities of LLMs, leading to more accurate and interpretable recommendations. Our key contributions can be summarized as follows:
@@
  • To the best of our knowledge, this work is among the first to explore the application of the graph of thoughts framework in the context of sequential recommendation.
@@
  • We introduce GOT4Rec, which optimally leverages the reasoning capabilities of LLMs to comprehensively capture and utilize the information within user sequences.
@@
  • Extensive experiments conducted on three datasets demonstrate that our method outperforms neural and LLM-based sequential models as well as other LLM reasoning strategies.
@@ 2 Related Work
2.1 Sequential Recommenders
@@
Traditional neural sequential recommendation systems aim to capture sequential dependencies in user behavior sequences to model dynamic user preferences, as seen in methods like GRU4Rec Hidasi et al. (2016). Techniques like self-attention and graph neural networks have further advanced the field Kang and McAuley (2018); Zhou et al. (2020); Wu et al. (2019); Zhang et al. (2022, 2023). These methods however, predominantly rely on sequence modeling capabilities and often inadequately incorporate textual information. Recently, research on transferable item representations Hou et al. (2022); Li et al. (2023a) has gained attention. These approaches remain constrained by limited datasets and fail to fully leverage real-world knowledge. With the advent of LLMs, new potential are offered by leveraging LLMs’ language understanding for recommendations. They can enhance features Xi et al. (2024); Lin et al. (2024); Liu et al. (2024) or act as rankers Dai et al. (2023); Harte et al. (2023); Li et al. (2023b); Sanner et al. (2023); Hou et al. (2024b). For example, ReLLa Lin et al. (2024) employs semantic user behavior retrieval to improve data quality and introduces retrieval-enhanced instruction tuning to enhance few-shot recommendation. LLMRank Hou et al. (2024b) formalizes the recommendation problem as a conditional ranking task and utilizes LLMs as zero-short rankers. Although these approaches are promising, existing research has not yet fully capitalized on the reasoning capabilities of LLMs.
2.2 LLM Reasoning Strategies
@@
Numerous approaches have been proposed to exploit the reasoning capabilities of LLMs. Chain-of-thought (CoT) Wei et al. (2022) introduces intermediate reasoning steps to enhance LLM performance, while Chain of Thought with Self-Consistency (CoT-SC) Wang et al. (2023) refines this by generating multiple CoTs and selecting the best output. Tree of thoughts (ToT) Yao et al. (2024) and graph of thoughts (GoT) Besta et al. (2024) further extend these methods by modeling reasoning as a tree or graph, respectively, to better generate and aggregate different thoughts, thereby improving complex tasks. In the recommendation domain, SLIM Wang et al. (2024) utilizes a CoT-based knowledge distillation module to transfer the step-by-step reasoning capabilities from a teacher model to a student model. However, its reliance on basic CoT limits its ability to fully capture the abundant sequential dependencies and intricate patterns within user interactions.
@@
Figure 2: An example of graph decomposition in the proposed GOT4Rec method. The user interaction sequence is divided to facilitate different types of reasoning: the last few items are used for short-term preference reasoning, while the entire sequence informs long-term preference reasoning. Additionally, sequences from similar users are retrieved for collaborative reasoning. In the thought graph, these thoughts are generated and subsequently aggregated to capture and integrate the various aspects of information in the user interaction sequence.
@@ 3.1 Overview of Thought Graph
When interacting with LLMs, we input messages (prompts) and LLMs respond with generated outputs (thoughts). Building on the GoT framework Besta et al. (2024), we model our GOT4Rec method as a tuple $(G,T)$, where $G$ represents the reasoning process and $T$ denotes the thought transformations. Specifically, we model the reasoning process as a directed graph $G=(V,E)$, where $V$ is the set of vertices and $E\subseteq V\times V$ is the set of edges. Each vertex represents a thought, encapsulating a solution to the current recommendation step along with other relevant global information. A directed edge $(t_{1},t_{2})$ signifies a dependency between thoughts, indicating that thought $t_{2}$ is generated by LLMs based on $t_{1}$.
In our method, we employ the generation transformation $T_{G}$ and aggregation transformation $T_{A}$ from GoT framework. The generation transformation generates one or more new thoughts based on an existing thought $v$. In this operation, new vertices and edges are generated: $V_{+}=\{v_{1+},...,v_{k+}\}$ and $E_{+}=\{(v,v_{1+}),...,(v,v_{k+})\}$, where $v_{1+},...,v_{k+}$ are new thoughts generated by $v$. Reasoning steps such as CoT can be incorporated into this process. While the aggregation transformation combines multiple thoughts into a consolidated thought, reinforcing their strengths while mitigating their weaknesses. A new vertex $v_{+}$ is created in this operation: $V_{+}=\{v_{+}\}$ and $E_{+}=\{(v_{1},v_{+}),...,(v_{k},v_{+})\}$, where $v_{1},...,v_{k}$ are the aggregated thoughts.
@@
3.2 Short-Term & Long-Term Reasoning Process
Numerous studies have emphasized the importance of understanding both a user’s dynamic short-term and stable long-term preferences in recommendation systems Yu et al. (2019); Zheng et al. (2022). In our GOT4Rec method, we enable LLMs to extract these two types of preferences separately and then synergize the most relevant aspects of each. To reason about short-term preferences, we select the last few interactions from the user sequence $S_{u}$ as $S_{\text{short}}$. We then apply and enhance the zero-shot CoT strategy from Wang et al. (2024) within the generation transformation $T_{G}$ to effectively capture the user’s short-term preferences and identify three categories that the user is likely to favor. As illustrated in Figure 3 as summarizing prompt, the generation transformation $T_{G}$ involves two key steps:
  • Step 1. Summarize user’s short-term preferences based on the given $S_{\text{short}}$.
@@ 3.2 Short-Term & Long-Term Reasoning Process
For each identified category, we prompt the LLMs to generate $N$ items the user is most likely interested in, repeating this process three times. The prompt template is provided in Figure 3 as recommendation prompt. After generating the item sets, we utilize the aggregation transformation $T_{A}$, where the three item sets undergo a voting process by the LLMs. This results in a new set of $N$ items that are most likely to interest the user within each category. After aggergation, we have three final item sets, which are subjected to a final round of voting by the LLMs to determine the top-$N$ items $I_{\text{short}}$ that best represent the user’s short-term preferences. The voting prompt is also provided in Figure 3.
To capture long-term preferences, we extend the input to include the entire user interaction sequence. Similar to the short-term reasoning process, we employ the zero-short CoT strategy within the generation transformation $T_{G}$, as shown in Figure 3. This allows LLMs to effectively differentiate and identify short-term and long-term preferences. After generating the relevant items, the final set of top-$N$ items $I_{\text{long}}$, representing the user’s long-term preferences, is determined through the aggregation transformation $T_{A}$.
@@ 3.3 Collaborative Reasoning Process
Collaborative filtering is widely used in various recommendation scenarios, modeling users’ preference based on their past interactions. In our method, we leverage the all-mpnet-base-v2 Reimers and Gurevych (2020), a sentence-transformers model that maps sentences to a vector space, to generate an embedding vector for the current user’s interaction sequence, allowing us to retrieve a set of sequences $S_{co}$ from other users most similar to the current user’s sequence. Details of this retrieval process are provided in the following section.
Once we have obtained the similar user sequences $S_{\text{co}}$, we employ the zero-short CoT strategy to generate three sets of top-$N$ items that the current user may be interested in. The two-step generation prompts, illustrated in Figure 3 as collaboration prompt, are as follows:
  • Step 1. Summarize the shared preferences between the current user and other users based on the given $S_{\text{co}}$.
  • Step 2. Based on the summarized preferences from Step 1, recommend $N$ items selected from $S_{\text{co}}$ that the current user is likely to prefer.
The three sets of items generated through this process are then aggregated and selected by the LLMs using the aggregation transformation $T_{A}$. This results in the final set of top-$N$ items $I_{\text{co}}$ that best reflect the collaborative preferences of users with similar interests to the current user.
At this point, we have obtained three distinct sets of items: $I_{\text{short}}$, which reflects the user’s short-term preferences; $I_{\text{long}}$, which captures the user’s long-term preferences; and $I_{\text{co}}$, which represents the collaborative preferences derived from other users. Finally, using the aggregation transformation $T_{A}$, we enable LLMs to synthesize these multiple sources of information to generate the final set of top-$N$ recommended items $I_{\text{fin}}$. The prompt template for this final aggregation is the same as the voting prompt.